40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 314 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  1167   Tue Dec 2 19:18:10 2008 ranaSummaryPEMRanger SS-1
In entry http://dziban.ligo.caltech.edu:40/40m/881 and a follow up from Jenne I put in the Ranger calibration.
Since then, we've reduced the SR560 gain from 200 to 100 so the calibration factor is now:

1e-9 (m/s)/count and then 2 poles at 0 Hz, and a Q~1 zero pair at 1 Hz.
in DTT:
G = 1e-9
p = 0, 0
z = 0.7 0.7
  1166   Tue Dec 2 17:56:56 2008 Alberto, RanaConfigurationPSLMC Alignment
In the attempt to maximize the Mode Cleaner transmission and minimize the reflection from the steering mirrors of the MC periscope, we could not get more ~2 V at the MC Trans PD and ~ 0.5 V at MC REFL_DC. As it turned out from the SUS Drift Monitor, the reason was that the MC optics had been somehow displaced from the optimal position.

After restoring the reference position values for the mirrors and tweaking again the periscope, we got ~3V at the MC TransPD and 0.5V at the reflection.
The beam was then probably clipped at the REFL PD so that we had to adjust the alignment of one of the BS in the transmitted beam path on the AS table.
We also zeroed the WFS PDs, but not before reducing the power from the MZ, for their QPDs not to saturate.

After relocking, the transmission was 3V and the reflection ~0.3V.

The beam isnow centered on the Trans PD and REFL PD and the Mode Cleaner locked. More details on the procedure will follow.
  1165   Mon Dec 1 15:09:27 2008 robUpdatePEMhalf-micron particle count is alarming
  1164   Thu Nov 27 22:56:42 2008 ranaConfigurationEnvironmenttemperature
8-)
Attachment 1: mc.png
mc.png
  1163   Tue Nov 25 19:29:15 2008 rana, alberto, johnConfigurationEnvironmenttemperature

Quote:
The PSL Room Temperature was alarming because it had gone above 23 C. This set off an unfortunate chain of events:

We found that the PSL HEPA was set low (20%). This is a fine setting for when no one is working in there but it
does raise the temperature since there are heat sources inside the blue box.

We tried to change the office area temperature to compensate and also the westmost sensor inside the lab area by 2 deg F.

The office area one was problematic - there was so much dust in it that the gas valve nipple was clogged. So we've
now blown it all clean with a compressed air can. We're now tuning the calibration screw to make our new
digital sensor agree with the setpoint on the controller.

Expect wild temperature swings of the office area for a couple days while Alberto and I tune the servo.


This morning Bob found 92F in the office area and in the control room of the lab. He turned down the thermostat and when I got in at about 9 I found 65. After a few hours of adjustment of the thermostat's calibration, I could stabilize the room temperature to about 72F. I also turned down the thermostats inside the lab of a couple of degrees F.
  1162   Tue Nov 25 18:38:03 2008 Alberto, RobUpdatePSLMC Periscope Alignment
This morning when I came in I found the MC cleaner unlocked and the autolocker script could not lock it. The reflected beam was quite off and showed in the bottom left corner of the IMCR camera. After turning off the WFS locking, I started slightly changing the alignment of the steering mirrors on the MC periscope, waiting for the LSC servo to lock the cavity. It didn't work. At some point I lost the beam from the IMCR camera and that is how someone might have found it when I left it for about one hour.

When I came back and tried again adjusting the steering mirrors, I noticed that the autolocker was working and was trying to lock the cavity. After just a bit of adjustment, the MC got easily locked.

After that, I spent a couple of hours trying to improve the alignment of the periscope to minimize the reflection and maximize the transmission. I started with a transmission of 0.4 V but, despite all the tweaking (I used the technique of turning both yaw knobs at the same time), I couldn't get more than 1.2 V (and 2.4 V at the reflection) if only the LSC servo was on. Looking at the camera, I moved the beam around to look for a more favorable spot but the MC wouldn't lock with the beam in other places. Maybe I could do better or maybe not because the cavity is not aligned. I'm going to try again tomorrow.
  1161   Mon Nov 24 19:15:16 2008 rana, alberto, johnConfigurationEnvironmenttemperature
The PSL Room Temperature was alarming because it had gone above 23 C. This set off an unfortunate chain of events:

We found that the PSL HEPA was set low (20%). This is a fine setting for when no one is working in there but it
does raise the temperature since there are heat sources inside the blue box.

We tried to change the office area temperature to compensate and also the westmost sensor inside the lab area by 2 deg F.

The office area one was problematic - there was so much dust in it that the gas valve nipple was clogged. So we've
now blown it all clean with a compressed air can. We're now tuning the calibration screw to make our new
digital sensor agree with the setpoint on the controller.

Expect wild temperature swings of the office area for a couple days while Alberto and I tune the servo.
  1160   Mon Nov 24 17:14:44 2008 ranaUpdatePSLMach Zender trends
It looks like the MZ has gotten less drifty after the alignment on Friday.
Attachment 1: Untitled.png
Untitled.png
  1159   Mon Nov 24 16:43:34 2008 ranaConfigurationComputersAlex and Jay took away some computers from the racks
I was over at Wilson house and saw Jay and Alex bring in 3 rackmount computers. One was a Sun 4600 and
then there were 2 3U black boxes. I got the impression that these were the data concentrators or
data collectors or framebuilder test boxes. They said that they got these from the 40m and no one was
in the lab to oppose them except for Bob and he didn't put up much of a fight.

Everything looks green on the DAQ Detail and RFM network screens so perhaps everything is OK. Beware.
  1158   Sat Nov 22 10:55:51 2008 CarynConfigurationIOODrum modes Lock-In settings changed
I unhooked the MC Demod Board's Qmon signal from the Lock-In. Set the demodulation frequency to 31.11Hz with 1V amplitude, and
put the output into MC_DRUM1. DTT showed a ~30Hz peak. Dataviewer showed signal with amplitude ~20,000.
Otherwise the settings were as Rana had them: Time Constant-100us,24dB/Sensitivity-200us/Low Noise
Want to check if Lock-In frequency drifts.
  1157   Fri Nov 21 21:28:32 2008 ranaSummaryComputersc0daqawg restart
A few minutes after restarting fb0 for the Guralp channels, the DAQAWG lights went red on the DAQ screens.
Why?? I chose revival procedure #3 for c0daqawg from the Wiki and it came back in a couple minutes.
  1156   Fri Nov 21 21:20:24 2008 ranaUpdatePEMGuralp noise measurement
This is the spectra and time series of the Guralp channels along with the Ranger (MC2). Looks like we could reduce the gain
on the ranger. The Guralp channels run into ADC noise around 40 Hz (which is OK). We'll have to look at the weekday trends
to see if they saturate.
Attachment 1: a.png
a.png
  1155   Fri Nov 21 20:29:43 2008 rana, alberto, robUpdatePSLMach Zender alignment

Quote:
So now the MZ alignment seems good; Alberto is on the MC periscope alignment like a cheap suit.


And alignment is now mostly done - MC locks on the TEM00.
                 REFL_DC

Unlocked           4.50  V
Locked noWFS       1.30  V
Locked + WFS       0.42  V
and the 29.5 MHz modulation depth is really small.

We should be able to rerun the Wiener analysis on this weekends MC data.

I don't know what our nominal StochMon numbers now are, but after Alberto tweaks up the alignment he can tell us if the RFAM has gotten any better.
  1154   Fri Nov 21 19:47:26 2008 ranaUpdatePEMGuralp noise measurement
and here's the spectra with them connected - from the coherences, it looks like it needs to be rotated by 90 deg.

I'll next rename the channels to fix this so that we get good seismic data over the weekend with the MC.
Attachment 1: a.png
a.png
  1153   Fri Nov 21 17:27:47 2008 JenneUpdatePEMGuralp noise measurement
Here is the data from the Guralp Seismometer for the past day or so, before I fixed the VERT-1 channel. The NS and EW show what's going on in the world, and VERT is measuring essentially the noise of the box, through the ADC, in counts.
Attachment 1: guralp_vert_shorted.jpg
guralp_vert_shorted.jpg
  1152   Fri Nov 21 16:52:48 2008 JenneUpdatePEMGuralp seismometer's Channel Problems are solved
PROBLEM noticed earlier this week: It looked like one of the seismometer channels (VERT-1) wasn't working, no matter how I put which channel into which input of the PEM ADCU. Watching the channel on Dataviewer, it looked like the ADC was measuring VERT-1 to be zero (actually measuring zero, not digital noise-type zero). I had checked the ADC by putting in a sine wave with a function generator, and saw on Dataviewer the wave I expected, so I knew that I had the correct channel, and that the channel was good.

SOLUTION: This afternoon I took the box out of the rack and opened it up. As soon as I opened it, I saw that I had left something inside the box which was causing the problem. Back when we were measuring the noise of the box, to ensure that it is lower than the ADC's noise, Rana and I had shorted the test points on the input of the VERT-1 channel with a little piece of wire. It turns out that I had closed up the box without remembering to remove the wire.

CONCLUSION of the story: I took out the piece of wire, and now all three seismometer channels (VERT-1, N/S-1, E/W-1) all work, and all detect me jumping around near the BSC. Since the seismometer breakout box reads a differential measurement, and since the input test points were connected, it was indeed measuring zero. Zero equals zero is all well and good, but it's even better now that it's measuring actual seismic motion.
  1151   Fri Nov 21 16:11:26 2008 rana, alberto, robUpdatePSLMach Zender alignment
The Mach Zender's dark port DC voltage had gone up too high (~0.5 V)and was turning yellow
on the screen. I re-aligned by touching the knobs on the 166 MHz path. Doing alignment after the
166 EOM had very little effect. The main improvement came from doing yaw on the turning mirror
just ahead of the 166 MHz EOM; this is the one which as no adjustment knobs (duh).

During this procedure, I had the MC off, the ISS off, and the MC autolocker off. After finishing
the alignment, the power on the ISS PDs had railed and the dark port power was ~0.29 V. So we
put in a ND0.2 on the ISS path and now the voltages are OK (~2 V on each PD). We have to remove the
ND filters and change the first ISS turning mirror into a ~10-20% reflector.


So now the MZ alignment seems good; Alberto is on the MC periscope alignment like a cheap suit.
Attachment 1: PB210051-1.JPG
PB210051-1.JPG
  1150   Fri Nov 21 16:03:50 2008 ranaHowToGeneralRecharging Batteries
I found some black & red "Ninja" alkaline AA batteries in the battery charger. This is dangerous. Please
do not put alkaline batteries in there, only Nimh. If you need help with the battery charger you can
come and talk to me or Rob and we can help you out getting started.
  1149   Thu Nov 20 09:37:39 2008 steveUpdatePSLMZ vs temp
This 12 days plot shows that it can hold lock if the daily temp variation is not more than 3/4 of a degree C
The MZ is happy if it's HV is above 100V
Attachment 1: mztemp.jpg
mztemp.jpg
  1148   Wed Nov 19 18:12:35 2008 ranaConfigurationIOOnew channel for MC drum modes
I set up the lockin to take the MC Demod Board's Qmon signal, demodulate it at 27.5 kHz, and
put the output into a DAQ channel (I think its either MC_DRUM1 or MC1_TEMPS). However,
the MC_DRUM channel doesn't look like its getting anything in the DTT although it looked fine
on a scope. I used the 'sensitivity' setting of the lockin to make the demodulated signal
large enough but not so large that it would saturate the ADC (+/- 2V).
  1147   Wed Nov 19 18:02:18 2008 ranaConfigurationElectronicsAll the Marconi Set to the Rubidium Frequency Standard
Not sure what was going on before. I changed the frequency counter to use an AC coupled input, and had it average
for 50 seconds. The output now agrees with the Marconi front panel to less than 1 Hz. Its still not 0.000 Hz,
but not bad.
  1146   Wed Nov 19 10:32:11 2008 AlbertoConfigurationElectronicsAll the Marconi Set to the Rubidium Frequency Standard
I placed the SRS Rubidium FS275 over the PSL rack, next to the frequency counter. This one and the Marconi on the PSL rack have been connected to the 10MHz output of the frequency standard. I set also the first Marconi, the one that used to drive the others, to external, direct frequency reference. Now it reads 166981718 Hz versus 166981725 Hz measured by the frequency counter: 8 Hz difference.
  1145   Tue Nov 18 19:44:53 2008 AlbertoUpdateGeneralX Arm Cavity "Negative" FSRs Measured
Previous measurements on the X arm cavity revealed a shift of the frequencies of the cavity resonances from where one would expect these to be by just looking at integer multiples of the cavity FSR. In particular, plotting the resonant frequencies versus the order of their occurrences while sweeping the laser frequency (in our case that of the beat between the two lasers), the linear fit of the data contained an unwanted offset:

resonant_frequency = n x FSR + offset

In part, we attributed this offset to the local oscillator of the PLL, the Marconi, which was not referred to an absolute frequency clock.
For that reason, I connected the Marconi to the RS FS275 which uses the 1PPS from the GPS to generate a 10 MHZ reference signal, and then scanned the cavity again. This time I started from negative beat frequencies, that happen when the frequency of the secondary laser is smaller than the main laser's, to positive frequencies. The way I made sure of the sign of the frequency was looking at the effect of changing the temperature of the NPRO. I decided that negative frequencies where those for which an increase in temperature lowered the beat frequency and positive frequencies those for which increasing the temperature made the beat frequency go up.
I then plotted the data and obtained the attached plot.

The offset was reduced to about 80 Hz (from more than 200 in the previous measurements). I think the residual offset has to do with something that happens in the cavity, something, as Koji found out, related to the alignment of the mirrors.

Thanks to the more data points, the measurement of the FSR improved to (3897627 +/- 5) Hz, which would let us know the measure of the cavity length with an error of 50um, if it weren't for the offset. I have to understand whether and how to take this into account to determine the precision in the cavity length. I guess it depends on whether it is real or it is still a systematic error due to the measurements.
Attachment 1: 2008-11-17_Linear_Fit.pdf
2008-11-17_Linear_Fit.pdf
  1144   Tue Nov 18 19:37:23 2008 YoichiUpdateIOOMC1 OSEM signals sign flipped and c1susvme1 restart problem
Around 2PM, MC1 started to swing crazily.
The damping feedback was not working and it was actually exciting the mirror wildly.
It turned out that the sign of the UR and UL OSEM signals flipped at that time.
Restarting c1sosvme fixed the problem.

While I was looking for the cause of the problem, c1susvme1 and c1susvme2 failed several times.
I don't know if it is related to this problem.
Now it is not trivial to restart c1susvme1. It fails to restart if you just power cycle it.
Alberto and I had to connect an LCD and a keyboard to it to see what was going on. After pushing the reset button on the front panel,
I had to press Ctrl+x. Otherwise, the state LED of c1susvme1 stays red and nothing happens.
After Ctrl+x, the boot screen came up but the boot sequence failed and an error message something the following was shown:
"PXE Boot failed, check the cable"
So I swapped the network cable with c1susvme2, which was already up and running.
This time, c1susvme1 started fine and surprisingly, c1susvme2 stayed alive.
Currently, both c1susvme1 and c1susvme2 are up and running with the LAN cables swapped.
We have to check the LAN cables.
  1143   Tue Nov 18 13:28:08 2008 CarynDAQIOOnew channel for MC drum modes
Alberto has added a channel for the Mode Cleaner drum modes.
C1:IOO-MC_DRUM1
sample rate-2048
chnum-13648
  1142   Mon Nov 17 20:47:19 2008 CarynSummaryGeneralDrove MC at 28kHz to excite drum modes
Rana, Alberto and I observed drum mode frequencies at 23.221kHz(MC1), 28.039kHz(MC2), 28.222kHz(MC3) while driving the mode cleaner. We observed no peaks when we didn't drive the mode cleaner. We used the SR785 to send a ~80mV noise signal in the 28-28.2kHz band to the mode cleaner mirrors via 1Y4-MC1,2,3-POSIN. Then we looked at 1Y2-Mode Cleaner-Qmon on the SR785 and saw peaks.
  1141   Mon Nov 17 16:59:22 2008 JenneUpdatePEMSeismometer hooked up, reading channels on DataViewer
Alberto, Jenne

The Guralp Seismometer is (finally) hooked up to the PEM ADCU. Alberto helped me make channels in the c1pem1 .ini file, which correspond to:

Guralp1 VERT = channel 9 on PEM ADCU = C1:PEM-SEIS_MC1_VERT
Guralp1 NS = channel 10 on PEM ADCU = C1:PEM-SEIS_MC1_NS
Guralp1 EW = channel 11 on PEM ADCU = C1:PEM-SEIS_MC1_EW

We also renamed the Ranger seismometer's channel to C1:PEM-SEIS_MC2_Y from C1:PEM-SEIS_MC1_Y, since tomorrow I'll move the Ranger Seismometer to be underneath MC2's chamber (it's currently sitting somewhere in the middle of the Mode Cleaner).


I can see the VERT and NS channels with dataviewer, but EW looks dead. I need to figure out if this is a bad cable thing, or if the ADC channel is no good, or if something in the box on that channel is no good. All 3 channels were tested and working after all the soldering was completed by Caryn, but something may have come undone while putting the box into its new place in the top of 1Y1. (In dataviewer, it looks like the EW channel is just floating, and not connected to anything.)

Plan of Attack:
* figure out why EW looks dead on Dataviewer
* redo Rana's static Wiener filter analysis, now that we have 2 seismometers (1 Ranger and 1 Guralp)
* work on adaptive Wiener filtering with the Guralp
  1140   Mon Nov 17 15:07:06 2008 YoichiUpdatePSLReference cavity ring down
I used MATLAB's system identification tool box to estimate the response of the reference cavity, i.e. cavity pole.
What I did was basically to estimate a model of the RC using the time series of the measured input and output power.

First, I prepared the input and output time series for model estimation.
The input is the input power to the RC, which I produced by inverting the PBS reflected light power and adding an offset
so that the signal is zero at t=0. Offset removal was necessary to make sure that the input time series does not give an
unintentional step at t=0.
The output time series is the transmission power of the RC. I also added an offset to make it zero at t=0.
Then I commanded MATLAB to compute the response of a first order low-pass filter to the input and try to fit
the computed response to the measured output by iteratively changing the gain and the cut-off frequency.
("pem" is the name of the command to use if you are interested in).

The result is shown in the attachment.
Blue curve is the input signal (I added a vertical offset to show it separately from the output).
The green curve is the measured output (RC transmission). The red curve is the response of the estimated model.
The estimated cut-off frequency was about 45kHz.

You can see that the red curve deviates a lot from the green curve after t=15usec.
By looking at this, I realized that the bandwidth of the RC cavity servo was too high.
The time scale we are looking at is about 50kHz whereas the FSS bandwidth is about 400kHz.
So when the input light was cut off, the error signal of the FSS becomes meaning less and the
input laser frequency was quickly moved away from the resonance. This is why the green curve does not
respond to the large peaks in the blue curve (input). The cavity was already off-resonance when the input power
showed bumps.

Since the red curve matches nicely with the green curve at the very beginning of the ring down, the estimated 45kHz
cavity pole is probably not that a bad estimate.

To make a better measurement, I will try to reduce the bandwidth of the RC servo by using only the PZT actuator.
If there were no ringing in the input light power, we wouldn't have to worry about the bandwidth of the servo because our
feedback is all made to the laser, not the cavity length.
In order to reduce the ringing in the input power, I asked Bob to make new HV cables using HV grade coax cables.
Attachment 1: Fit.png
Fit.png
  1139   Mon Nov 17 11:01:15 2008 AlbertoHowToElectronicsCalibrating the Frequency Standard of the Marconi
I locked the SRS Rubidium Frequency Standard FS275 to the the 1pps from the GPS. The specs from the manual provide a frequency accuracy of 5x10^-11, that is 5x10-4 @ 10 MHz, since this is the reference signal frequency we're are going to use.
The Marconi internal frequency standard is provided by a TCXO oscillator. The instrument can be set in either one of these ways: 1) Indirect Synchronization, by which the internal TCXO is phase-locked to the external frequency standard (i.e. the SRS FS275 in our case) 2) Direct Sync, in which the internal TCXO is bypassed and the frequency standard is the external one.

I checked the specs of both frequency standards and found:

SRS FS275: 5x10^-11 -> 5x10^-10 Hz @ 10 MHz

Marconi: here what the data sheet says is that "the temperature coefficient is 7 in 10^7 in the temperature range between 0 and 55 C" and so should be also the frequency accuracy.

The SRS FS275 seems more accurate than the TCXO therefore I'm going to set the Marconi on the direct external mode.
Attachment 1: 2023ASeriesOperatingManual.pdf
2023ASeriesOperatingManual.pdf 2023ASeriesOperatingManual.pdf 2023ASeriesOperatingManual.pdf 2023ASeriesOperatingManual.pdf
Attachment 2: SRS_FS275_Rubidium_Frequency_Standard.pdf
SRS_FS275_Rubidium_Frequency_Standard.pdf SRS_FS275_Rubidium_Frequency_Standard.pdf
  1138   Fri Nov 14 22:40:51 2008 YoichiUpdatePSLReference cavity ring down

Quote:

To make the DEI pulser make a fast pulse on the EO shutter EOMs, we had to make sure:

1) the cable had a high voltage rated dielectric. cheap dielectrics show the 'corona'
effect, especially when there is a bend in the cable.


I'll check it with Bob.


Quote:

2) the EO has to have a resistor on it to prevent ringing due to the impedance mismatch.


Did you use a shunt or series resistor ?
If shunt, I guess it has to have a huge heat sink.
Actually, DEI says the pulser does not require any external shunt/series resistors or impedance-matching network.
Looks like it is not true ...


Quote:

3) We needed ~3.5 kV to get the EO shutter crystal to flip the light by 90 deg.


Yes, I adjusted the voltage to maximize the power change and it was about 3.5kV.
  1137   Fri Nov 14 20:35:47 2008 ranaUpdatePSLReference cavity ring down
To make the DEI pulser make a fast pulse on the EO shutter EOMs, we had to make sure:

1) the cable had a high voltage rated dielectric. cheap dielectrics show the 'corona'
effect, especially when there is a bend in the cable.

2) the EO has to have a resistor on it to prevent ringing due to the impedance mismatch.

3) We needed ~3.5 kV to get the EO shutter crystal to flip the light by 90 deg.
  1136   Fri Nov 14 19:20:42 2008 YoichiUpdatePSLReference cavity ring down
Thanks to Bob making the high-voltage BNC cables for the HV pulse generator, I was able to operate the EOM in front of
the reference cavity.

The conceptual setup is the following:
[HV pulse] ----+           +-->-- [PD2]
               V           |
->--[HWP]->-- [EOM] -->-- [PBS] --<->-- [QWP] --<->-- [Reference Cavity] -->-- [PD1]
                           |
                [PD3] --<--+

The high voltage pulse rotates the polarization of the light after the EOM. When the HV is applied, the PBS reflects most of the light
into PD2 (Thorlabs PDA255), shutting down the incident light into the cavity.
The transmitted light power of the reference cavity is monitored by PD1 (PDA255). The reflected light from the reference cavity
is monitored by the DC output of the RF PD (PD3). PD3 is low-passed so the response is not fast.
Thorlabs says PDA255 has 50MHz bandwidth.

The attached plot shows the time series of the above PD signals when the HV was applied.
Input Pulse (blue curve) is the input to the HV pulse generator. When it is high, the HV is applied.
"PBS reflection" (red) is PD2. "Reflection" (green) is PD3. "Transmission" (light blue) is PD1.

The red curve shows huge ringing. At first I thought this was caused by the bad response of the PD.
However, the same ringing can be seen in the PD3 and the peaks match very well.
When red curve goes down the green curve goes up, which is consistent with the energy conservation.
So it looks like the light power is actually exhibiting this ringing.
May be the HV pulse is distorted and the voltage across the EOM is showing this ringing.
I will check the input voltage shape to the EOM using a high impedance probe, if possible.

The green curve shows a slow decay because it has a long time constant. It is not an actual
trend of the reflected light power.

The RC transmission power shows some peaks, probably due to the ringing in the input power.
So just fitting with an exponential would not give a good estimate of the cavity pole.
Even though, we should be able to de-convolute the frequency response of the reference cavity
from the input (red curve) and output (light blue curve) signals.
Attachment 1: RingDown.png
RingDown.png
  1135   Fri Nov 14 17:41:50 2008 JenneOmnistructureElectronicsSweet New Soldering Iron
The fancy new Weller Soldering Iron is now hooked up on the electronics bench.

Accessories for it are in the blue twirly cabinet (spare tips of different types, CD, and USB cable to connect it to a computer, should we ever decide to do so.

Rana: the soldering iron has a USB port?
Attachment 1: newSolderingIron.JPG
newSolderingIron.JPG
  1134   Fri Nov 14 11:33:19 2008 steveUpdateVACSRS-RGA installed
Old Dycor rga is removed and new SRS-RGA200 installed.
It is pumped down and ready for hooking up its RS-232 output for operation.
  1133   Thu Nov 13 15:50:44 2008 AlbertoConfigurationGeneralGPS 10MHz clock
The schematic of the 1Y7 rack that Alan pointed out (see attachment) don't represent anymore the actual rack.
First, with Yoichi we found that the GPS receiver for the 10 MHz is in a different position,
on the other side of the VME computer. It seems to output 1 kHz, which also appears in some modulated way.
This signal is then passed to a board on 1Y7 that seems make just copies of the signal. One of these goes
to an other board in 1Y6 that looks like a GPS receiver but has actually no GPs antenna input. Here it is
not connected to anything, but on its same crate is a the awg computer, so it might be providing the clock
to that by the crate.

We also checked the clock monitor output on the boards in the PSL that provides for the clock to the Penteks
and it seems that these are actually getting 4MHz.
Attachment 1: rackstuff.pdf
rackstuff.pdf rackstuff.pdf rackstuff.pdf rackstuff.pdf
  1132   Thu Nov 13 11:33:25 2008 AlbertoHowToTreasureMaking (good) Matlab figures
Just a little summary of some useful ways to change plot settings in Matlab that I wanted to share and remember for the future:

http://saig.phys.ualberta.ca/toolbox/Matlab/making_figures.html
  1131   Wed Nov 12 11:36:13 2008 JenneUpdatePEMGuralp Breakout Box is ~50% hooked up
The Guralp box is about halfway hooked up now. The seismometer is under the BSC, and the long cable from the seismometer to the breakout box is connected to "Guralps 1 Input" on the front panel. This corresponds to the set of 3 channels that Caryn stuffed with the new fancy-pants resistors few weeks ago. (When we finally get the other Guralp back from the company, we'll have to stuff the next set of 3 channels).

The Breakout Box is on the very top of 1Y1, sitting on top of the black power supplies. This should be fine, but it's pretty toasty hot up there, so if we find that there are problems with running the box at higher-than-room-temperature, step 1 will be to find a new spot for the box. (I'm not at this time anticipating a problem, but you never know....) Steve put a little foot between the Guralp box and the power supply to get some air circulation.

The ADC Octopus cable that Bob made is connected, and going up through the top of the rack. I am now going on a BNC cable hunt to extend this cable over to the PEM ADC. The PEM ADC is in 1Y7, so I'll need some medium-long BNC cable to get there.

The power cable is also ready to be connected to the rack's +/- 15VDC. I'll talk to Bob about getting this done.

Next step: pick some channels on the PEM ADC, and create them in the .ini files
  1130   Wed Nov 12 11:14:59 2008 CarynDAQPSLMC temp sensor hooked up incorrectly
MC Temperature sensor was not hooked up correctly. It turns out that for the 4 pin LEMO connections on the DAQ like J13, J14, etc. the channels correspond to horizontal pairs on the 4 pin LEMO. The connector we used for the temp sensor had vertical pairs connected to each BNC which resulted in both the differential pairs on J13 being read by the channel.
To check that a horizontal pair 4 pin LEMO2BNC connector actually worked correctly we unlocked the mode cleaner, and borrowed a connector that was hooked up to the MC servo (J8a). We applied a sine wave to each of the BNCs on the connector, checked the J13 signal and only one of the differential pairs on J13 was being read by the channel. So, horizontal pairs worked.
  1129   Mon Nov 10 19:39:40 2008 steveUpdatePEMthump noise from particle counter
The particle counter is back on the IOOC location on a piece of FOAM
It needs this isolation, so when the pump is running, it's not shaking things.
The counter was counting for 6 sec and it was on holding for 20 mins.
Now I set the counter for 20 sec so it is easy to recognise it's signal and it holds for 2min only.
This will set the alarm handler in action.

Atm1: 40 mins plot
PEM-ACC_MC2_x,y,z up to 13 mins: pcounter at MC2 table, clamped, counting for 20s and holds for 2 mins
PEM-ACC_MC2_x,y,z from 13 to 26 mins: pcounter at MC2 table, not clamped, seated on 2" foam, counting 20s and holds for 2 mins
PEM-ACC_MC1_x,y,z from 26 to 40 mins: pcounter at MC1_IOOC location, not clamped, seated on 2" foam, counting 20s and holds for 2 mins

Rana won the bet



Quote:
I'll bet Steve a dollar that it is mechanical. The attached PDF compares all of the accelerometers from right now.

You can see that the RMS in MC2 is way bigger than MC1.

In the second PDF file you can see the time series. I had to play around a lot with DTT to get it to work. The DTT/Foton
combo on Allegra is not stable, so make sure your work early and often.

In the plots shown, I am bandpassing the time series from 600-700 Hz. I found that doing so allowed the burp in MC1 to remain
large and reduce the extraneous fuzz in MC2. As you can see there is no such noise in MC2.

So its a noise around 600-700 Hz that comes on quickly and then shuts off after several seconds. Its also very periodic in that
it comes on around every 20 minutes. Steve also tells me (although he refuses to put in the elog) that it started up around
August 20th (?). I feel like someone in the 40m lab ought to be able to guess what this is at this point.

Please convince Steve to elog his findings about when the noise started.

If one goes out there and stands next to it when the trend predicts its happening it becomes clear what it is.
Attachment 1: partcfm.jpg
partcfm.jpg
  1127   Mon Nov 10 16:44:14 2008 steveUpdatePEMparticle counter at MC2
The particle counter was moved to MC2 table temporarily.
It is clamped to the table at the south west corner.
It is pumping for 20s and waiting for 2 minutes now.
  1126   Mon Nov 10 11:32:49 2008 robUpdateComputersc1iscex rebooted

it was running a few cycles late
  1125   Mon Nov 10 11:06:09 2008 robHowToIOOmode cleaner locked

I found the mode cleaner unlocked, with (at least) MC1 badly mis-aligned. After checking the coil alignment biases and finding everything there looking copasetic, I checked the trends of SUS{PIT,YAW,POS} and found that both MC1 and MC3 took a step this morning. The problem turned out to be loosed/jiggled cables at the satellite amplifiers for these suspensions. Giving them a good hard push to seat them restored the alignment and the mode cleaner locked right up.
  1124   Fri Nov 7 18:38:19 2008 AlbertoDAQPSLMC temperature sensor hooked up
Alberto, Rana,
we found that the computer handling the signals from ICS-110B was C1IOVME so we restarted it. We changed the name of the channel to C1:PEM_TEMPS and the number to 16349. We tracked it up to the J14 connector of the DAQ.
We also observed the strange thing that both of the differential pairs on J13 are read by the channle. Also, if you connect a 50 Ohm terminator to one of the pairs, the signal even get amplified.

(The name of the channel is PEM-MC1_TEMPS)
  1123   Fri Nov 7 16:05:55 2008 steveBureaucracySAFETYinsect killer sprayed at kitchen area !

Bob and I cleaned out the sink area and sprayed
Spectracide's BUG STOP insect killer solution on the shelfs and sink
table top area.

NO eating or coffee drinking till Monday

This is an effort to stop the ants coming.
  1122   Fri Nov 7 15:58:10 2008 ranaUpdatePEMAC is back on
I'll bet Steve a dollar that it is mechanical. The attached PDF compares all of the accelerometers from right now.

You can see that the RMS in MC2 is way bigger than MC1.

In the second PDF file you can see the time series. I had to play around a lot with DTT to get it to work. The DTT/Foton
combo on Allegra is not stable, so make sure your work early and often.

In the plots shown, I am bandpassing the time series from 600-700 Hz. I found that doing so allowed the burp in MC1 to remain
large and reduce the extraneous fuzz in MC2. As you can see there is no such noise in MC2.

So its a noise around 600-700 Hz that comes on quickly and then shuts off after several seconds. Its also very periodic in that
it comes on around every 20 minutes. Steve also tells me (although he refuses to put in the elog) that it started up around
August 20th (?). I feel like someone in the 40m lab ought to be able to guess what this is at this point.

Please convince Steve to elog his findings about when the noise started.

If one goes out there and stands next to it when the trend predicts its happening it becomes clear what it is.
Attachment 1: mc-acc.pdf
mc-acc.pdf
Attachment 2: mc-acc-quad.pdf
mc-acc-quad.pdf
  1121   Fri Nov 7 10:52:57 2008 steveUpdatePEMAC is back on
The 20 minutes and 6 sec thumps are not related to the 40m ac units
This period is precisely right on so it is not likely mechanical as a source.
The time and frequency domain of this signal at attachment 2&3

First I had the chilling water cut off and later I turned off the fans
as 5 hrs temp plot shows
Attachment 1: acf.jpg
acf.jpg
Attachment 2: acc_mc1.jpg
acc_mc1.jpg
Attachment 3: test.jpg
test.jpg
  1120   Fri Nov 7 08:08:00 2008 steveUpdatePEMAC turned off in IFO room
All tree AC units in IFO room 104 switched off at 7:45am
I'm looking for the periodic thump signal in the Vertex area
noted in entry 1113 of PEM-ACC_MC1
  1119   Thu Nov 6 22:07:56 2008 ranaConfigurationComputersELOG compile on Solaris
From the ELOG web pages:

Solaris:

Martin Huber reports that under Solaris 7 the following command line is needed to compile elog:

gcc -L/usr/lib/ -ldl -lresolv -lm -ldl -lnsl -lsocket elogd.c -o elogd

With some combinations of Solaris servers and client-side browsers there have also been problems with ELOG's keep-alive feature. In such a case you need to add the "-k" flag to the elogd command line to turn keep-alives off.
  1117   Thu Nov 6 10:06:41 2008 steveUpdateLockingarms lock degradation
I have been locking the arms in the mornings lately.
The daily drift of LSC-TRX is ~ 15% and LSC-TRY ~5%
Attachment 1: arms.jpg
arms.jpg
  1116   Thu Nov 6 09:45:27 2008 steveUpdateMOPAhead temp hick-up vs power
The control room AC temp was lowered from 74F to 70F around Oct 10
This hold the head temp rock solid 18.45C for ~30 days as it shows on this 40 days plot.
We just had our first head temp hick-up

note: the laser chiller did not produce any water during this period
Attachment 1: htpr.jpg
htpr.jpg
ELOG V3.1.3-